OpenStudy (lovelyharmonics):

Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by two.

OpenStudy (lovelyharmonics):

@Luigi0210

OpenStudy (luigi0210):

Do you know what co-terminal means?

OpenStudy (luigi0210):

*What co-terminal angles are

OpenStudy (lovelyharmonics):

no... and what is with your picture?

OpenStudy (luigi0210):

Pretty simple, just \(\pm 2\pi/360^○\) The pic is a long story, don't ask xD

OpenStudy (lovelyharmonics):

oh my so the answers will be +2 /pi /360 and -2 /pi /360?

OpenStudy (lovelyharmonics):

.-. darn i hate codes..... +2pi/360 and -2pi/360

OpenStudy (luigi0210):

Yup, so for the positive \(\large \frac{\pi}{2}+2\pi\) amd for the negative \(\large \frac{\pi}{2}-2\pi\)

OpenStudy (luigi0210):

Awe, it's alright ლ(゚д゚ლ)

OpenStudy (lovelyharmonics):

oh wait is it pi/2 + 360 and pi/2 -360

OpenStudy (luigi0210):

Simplify them silly girl >.> \(\large \frac{\pi}{2} \pm \frac{4\pi}{2}\)

OpenStudy (lovelyharmonics):

no it not .-. three pi divided by two; negative pi divided by two pi divided by two + 360°; pi divided by two - 360° five pi divided by two; negative pi divided by two five pi divided by two; negative three pi divided by two

OpenStudy (luigi0210):

Oh you >.< \[\LARGE \frac{\pi}{2}+\frac{4\pi}{2}=\frac{5\pi}{2}\] \[\LARGE \frac{\pi}{2}-\frac{4\pi}{2}=-\frac{3\pi}{2}\]

OpenStudy (lovelyharmonics):

wait but before you said +- 4pi/2

OpenStudy (luigi0210):

Yes, that's what you add/subtract to get the co-terminal angle.

OpenStudy (luigi0210):

Short explanation http://hotmath.com/hotmath_help/topics/coterminal-angles.html

OpenStudy (lovelyharmonics):

oh that makes sense so like 4pi-1pi

OpenStudy (luigi0210):

I just convert the \(2\pi\) into a fraction. It will either be \(\pm 2\pi/360^○\)

OpenStudy (lovelyharmonics):

but thats one of the answers too .-.

OpenStudy (luigi0210):

The answer is D >.<

OpenStudy (lovelyharmonics):

yay now unto more math c: